The following data shows frequency distribution of error page of a book. Fit Poisson distribution to the following :


Number error - 1,2,3,4,5,6,7,8


Frequency-12,18,19,05,09,14,12,11

1
Expert's answer
2021-12-30T05:57:48-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 x_i & f_i & x_if_i & P(X=x_i) & Expected\\
 & & & & value\\
\hline
 1 & 12 & 12 & 0.060161 & 6 \\
 \hdashline
 2 & 18 & 36 & 0.128143 & 13 \\
 \hdashline
 3 & 19 & 57 & 0.181963 & 18 \\
 \hdashline
 4 & 5 & 20 & 0.193791 & 19 \\
 \hdashline
 5 & 9 & 45 & 0.165110 & 17\\
 \hdashline
 6 & 14 & 84 & 0.117228 & 12 \\
 \hdashline
 7 & 12 & 84 & 0.071341 & 7 \\
 \hdashline
 8 & 11 & 88 & 0.037989 & 4 \\
 \hdashline
 9 & & & 0.044274 & 4 \\
 
 \hdashline
\end{array}"

"mean=\mu=\dfrac{\sum_ix_i f_i}{\sum_if_i}=\dfrac{426}{100}=4.26"

"\lambda=\mu=4.26"

"P(X=x_i)=\dfrac{e^{-\lambda}\cdot\lambda^{x_i}}{(x_i)!}"